1.   An investor is pondering on whether to invest in a siomai franchise or in a printing press company. He had his staff analyze the cost benefits that he would get from the two business ventures. Show the report of each business.

 

A.   The siomai franchise would need Php 800 000 as capital. It has a 60-percent chance of profiting Php 400 000; a 15-percent chance of profiting Php 600 000; and 25-percent chance of losing the capital entirely.

B.   The paper manufacturing company would need Php 800 000 as capital. It has a 35-percent chance of earning Php 1 000 000 and a 65-percent chance of losing the capital entirely.

C.  Which between the two business ventures should the investor invest his Php 800 000?



1
Expert's answer
2021-04-27T01:32:13-0400

Find the expected profit for case A:

"M(A) = 400000 \cdot 0.65 + 600000 \cdot 0.15 - 800000 \cdot 0.25 = 150000"

Find the expected profit for case B:

"M(B) = 1000000 \cdot 0.35 - 800000 \cdot 0.65 = - 170000"

Since in case A a profit is expected, and in case B - a loss, then the investor should invest in a franchise siomai (case A)